OpenStudy (anonymous):

A line has slope 1/4 and y-intercept 4. Which is the correct equation that describes this line in slope-intercept form?

OpenStudy (bahrom7893):

y = (1/4)x+4

OpenStudy (anonymous):

y=(1/4)x+4

Directrix (directrix):

y = mx + b y = ( 1/4 ) x + 4